# Adversarial Prompt Tuning for Vision-Language Models

### v2

@@ 5.1 Experimental Settings
5.2 Comparison with Vanilla CLIP
We started our evaluation by comparing AdvPT with the vanilla CLIP model. Using PGD-40 and RAP, we evaluated adversarial robustness in 8 datasets, as indicated in Tab. 1. Our findings reveal that: (1) AdvPT demonstrates improvements over the vanilla CLIP under both PGD-40 and RAP attacks, with the specific improvement quantified in green. Since RAP constitutes a black-box attack, the improvements are subtle. But under the white-box attack with PGD-40, there is a significant improvement. (2) While the primary goal of AdvPT is not to enhance generalizability, the empirical finding implys that the enhancement of accuracy emerges as a collateral advantage.
Refer to caption (a) Flowers
Refer to caption (b) Pets
Refer to caption (c) Food101
Refer to caption (d) SUN397
Refer to caption (e) DTD
Refer to caption (f) EuroSAT
Refer to caption (g) UCF101
Refer to caption (h) ImageNet
Figure 3: AdvPT vs. CoOp on generalization (clean accuracy) and adversarial robustness (robust accuracy).
5.3 Comparison with Adversarial Defenses
As previously described, AdvPT presents an innovative approach to enhance the robustness of image encoders against adversarial attacks by modifying only the textual input. This method is inherently synergistic with visual-modality input denoising defenses. We evaluated its performance against powerful white-box attacks, PGD-40, and observed its compatibility with defenses, as delineated in Tab. 2. Significantly, incorporation of AdvPT requires no specialized tuning for the purified images.
Remarkably, the synergy of AdvPT with baseline defense mechanisms sometimes yielded “$1+1>2$” contribution. For example, on the ViT-B/16 model applied to the Flowers dataset, AdvPT alone increases accuracy by 31.0% (from 6.40% to 37.40%), yet when combined with Super-resolution, it further improves the performance of Super-resolution by 46.60% (from 13.80% to 60.4%). These findings highlight the potential of this innovative synergy strategy to enhance adversarial defense by simultaneously modifying textual and visual inputs, warranting further investigation in future study.
5.4 Generalization-Robustness Trade-off
In this subsection, we delve into the effects of various learning objectives on the learnable vectors within prompt tuning. The primary goal of AdvPT is to bolster the adversarial robustness of the image modality in VLMs, particularly when fine-tuned on adversarial images. In contrast, we explore whether an objective like CoOp [44], which is fine-tuned on clean images for improved accuracy, affects adversarial robustness differently.
Our comparative analysis of AdvPT and CoOp unveils insights into their generalizability and adversarial robustness, as illustrated in Fig. 3. Our findings are twofold: (1) Intriguingly, AdvPT significantly outperforms CoOp in adversarial robustness, albeit at a slight cost to generalization. This highlights a potential trade-off between adversarial robustness and generalization in prompt tuning, aligning with conclusions drawn from traditional adversarial training [40]. (2) Although AdvPT sacrifices some generalizability, this drawback is mitigated as the model scale increases. Particularly on ViT-L/14, while also enhancing adversarial robustness, the narrowed generalizability gap makes AdvPT highly compatible with the ongoing trend towards scaling up models.
Table 4: The nearest words (distance value) for learnable vectors. N/A means non-Latin characters.
@@ 5.4 Generalization-Robustness Trade-off
5.5 Comparison with Adversarial Training
Refer to caption (a) ViT-B/16
Refer to caption (b) ViT-L/14
Figure 4: Efficiency Comparison between AdvPT and adversarial training on Pets.
In this subsection, we compare the efficiency of our method, which focuses on fine-tuning only the prompt, against traditional adversarial training. Specifically, we juxtapose AdvPT with PGD-10 adversarial training [18] on the Pets dataset, ensuring that both methods use an equivalent batch size. The comparative results are depicted in Fig. 4, including the time taken to compute the adversarial embedding bank $\mathbf{A}$ in the total time reported.
Our analysis reveals that AdvPT is significantly more time-efficient than traditional adversarial training, requiring at least an order of magnitude less time. Moreover, it demonstrates a superior enhancement in model performance, outperforming adversarial training by at least one order of magnitude in effectiveness. This efficiency advantage, exceeding 100$\times$, positions AdvPT as a notably superior solution for VLMs.
5.6 Evaluation on Domain Shift
A notable advantage of CLIP lies in its adaptability to domain shift. Thus, in this subsection, we evaluate the transferability of AdvPT in comparison to the vanilla CLIP in domain shift scenarios. The source dataset utilized is ImageNet, while the target datasets include ImageNetV2, ImageNet-Sketch, ImageNet-A, and ImageNet-R. The results presented in Tab. 3 elucidate that the proposed AdvPT outperforms the vanilla CLIP in terms of adversarial robustness, thereby validating its stability across varied domains.
5.7 Further Analysis
Figure 5: Effect of number of learnable vector on Pets.
Number of Learnable Vector
In Sec. 5.4, we observed similarities between adversarial prompt tuning and adversarial training. It is widely acknowledged within the adversarial training framework that a larger count of tunable parameters correlates with enhanced adversarial robustness. To discern whether this correlation persists within the context of adversarial prompt tuning, we conducted an empirical evaluation of its efficacy under different number of learnable vector $M\in[1,50]$, using the Pets dataset as an example. The empirical outcomes, as illustrated in Fig. 5, indicates while a minimal adversarial robustness is observed at $M=1$, this reduction is considered to be negligible. This suggests that the volume of tunable parameters does not constitute a constraint in adversarial prompt tuning. Instead, unlocking its potential efficacy warrants further investigation.
Interpreting the Learnable Vector
In this subsection, we aim to decode what the learnable vectors have captured. However, direct mapping of these learnable vectors to words is infeasible due to the optimization occurring within a continuous space, while word space is discrete. Therefore, we adopt a technique applied in the CoOp experiment, searching vocabulary for the nearest words to the learned vectors by Euclidean distance, as illustrated in Tab. 4. Contrary to the vectors learned in CoOp, which tend to correlate with specific datasets, the vectors from AdvPT exhibit a higher degree of non-associativity. These words are not intuitively understandable, exactly aligning with the non-robust features in adversarial images.
@@ 6 Limitations
First, the study’s evaluation of adversarial robustness is limited by the narrow range of attacks used. Traditional adversarial attacks, especially adaptive attacks, are not deployable directly in this context. Future research should focus on developing new attack methods specifically for VLMs to comprehensively assess their robustness. Second, the paper’s focus is restricted to image recognition tasks. Exploring the applicability of AdvPT to a broader array of tasks, such as Visual Question Answering (VQA) in advanced models like GPT-4V [22], is a worthwhile direction for future research. Lastly, visual prompts [13, 14] emerge as a promising research avenue, given their extensive trainable parameters, which could enhance adversarial robustness. Yet, it introduces additional branches to the model, thus falling into the model robustification category. Addressing the computational challenges posed by this approach remains an open question for further exploration.

### v3

@@ 5.7 Evaluation on Domain Shift
Figure 6: AdvPT vs. vanilla CLIP on distribution shift.
Figure 7: Effect of number of learnable vector on Pets.
A notable advantage of CLIP lies in its adaptability to domain shift. Thus, in this subsection, we evaluate the transferability of AdvPT in comparison to the vanilla CLIP in domain shift scenarios. The source dataset utilized is ImageNet, while the target datasets include ImageNetV2, ImageNet-Sketch, ImageNet-A, and ImageNet-R. The results presented in Fig. 7 elucidate that the proposed AdvPT outperforms the vanilla CLIP in terms of adversarial robustness, thereby validating its stability across varied domains.
5.8 Further Analysis
5.8.1 Number of Learnable Vector
In Sec. 5.4, we observed similarities between adversarial prompt tuning and AT. It is widely acknowledged within the AT framework that a larger count of tunable parameters correlates with enhanced adversarial robustness. To discern whether this correlation persists within AdvPT, we conducted an empirical evaluation of its efficacy under different numbers of learnable vector $M\in[1,50]$, using the Pets dataset as an example. The empirical results, as illustrated in Fig. 7, suggest that the volume of tunable parameters does not constitute a constraint in AdvPT. Instead, unlocking its potential efficacy warrants further investigation.
5.8.2 Interpreting the Learnable Vector
In this subsection, we aim to decode what the learnable vectors have captured. However, direct mapping of these learnable vectors to words is infeasible due to the optimization occurring within a continuous space, while word space is discrete. Therefore, we adopt a technique applied in the CoOp experiment, searching vocabulary for the nearest words to the learned vectors by Euclidean distance, as illustrated in Tab. 5. These words are not intuitively understandable, exactly aligning with the non-robust features in adversarial images.
Table 5: The nearest words for learnable vectors. N/A means non-Latin characters.
@@ 6 Limitations
First, the paper’s focus is restricted to image recognition tasks. Exploring the applicability of AdvPT to a broader array of tasks, such as Visual Question Answering (VQA) in advanced models like GPT-4V [23], is a worthwhile direction for future research. Second, visual prompts [13, 14] emerge as a promising research avenue, given their extensive trainable parameters, which could enhance adversarial robustness. Yet, it introduces additional branches to the model, thus falling into the model robustification category.
@@ 7 Conclusion and Discussion
Acknowledgements
This work was supported by National Key R&D Program of China (Grant No. 2022ZD0160103, 2023YFC3310700), National Natural Science Foundation of China (Grant No. 62172094, 62276067) and Science and Technology Commission of Shanghai Municipality (Grant No. 22511106102).